OpenStudy (anonymous):

5x-5=-8 would you subtract the inside and then multiply the inside?

rvc (rvc):

5x=-8+5

OpenStudy (anonymous):

its solve for x

rvc (rvc):

so do ahead

rvc (rvc):

what is -8+5?

rvc (rvc):

@math-geek

OpenStudy (mathstudent55):

You want to solve for x. That means you need x by itself on the left side. Two things are happening to x. It is being multiplied by 5, and a 5 is being subtracted from it. First you need to undo the subtraction of 5. To do that, you must add 5 to both sides of the equation. Then you need to undo the multiplication by 5. To undo that, you must divide both sides by 5.

OpenStudy (anonymous):

-3

rvc (rvc):

so x=-3/5

OpenStudy (anonymous):

options -2 -1 - 3 4 0

OpenStudy (anonymous):

3. -3/4

OpenStudy (ciarán95):

I presume you're looking to solve this equation for x.... The easiest way to do this is to first bring the '-5' over to the right hand side of the equation. In doing so, we change the sign, so '-5' becomes '+5'. So, 5x -5 = -8 becomes: 5x = -8 + 5 = -3 Now we need to get rid of the '5' from '5x'. If we divide both sides of the equation by 5, we get: 5x/5 = -3/5 The 5s on the left-hand side (above and below the fraction) will cancel, and we are left with 'x' on its own from which you can see the answer.

OpenStudy (anonymous):

answers 0